OpenStudy (anonymous):

Complete the equation of the graphed linear function in point-slope form. y - 6 =____ (x____)

OpenStudy (anonymous):

i think the one in () is 6

OpenStudy (anonymous):

@RANE

OpenStudy (anonymous):

@henrry123

OpenStudy (anonymous):

@lostsilverfangirl

OpenStudy (anonymous):

but i don't know the slope

OpenStudy (anonymous):

i would help you but i suck at slope sorry

OpenStudy (anonymous):

its kewl thanx tho :3

OpenStudy (anonymous):

im no good at slopes sorry

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

@Hotchellerae21

OpenStudy (anonymous):

ill try it tho

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

I know how to get the slope, but I'm not very familiar with point-slope form, give me just a minute

OpenStudy (anonymous):

i think the one in () is -6 :)

OpenStudy (anonymous):

mk

OpenStudy (sleepyjess):

You almost have x right. Point-slope form is \(\sf (y-y_1)=m(x-x_1)\) notice how it is subtracting x.

OpenStudy (anonymous):

oh

OpenStudy (sleepyjess):

So what is x-(-6) equal to?

OpenStudy (anonymous):

6x

OpenStudy (sleepyjess):

just 6

OpenStudy (sleepyjess):

So now we have \(\sf (y-6)=m(x+6)\). Now we need to find another point on that line to get the slope.

OpenStudy (sleepyjess):

Oh, it looks like there is a point at -8, 2. Are you familiar with the slope formula?